OpenStudy (anonymous):

Betty owns x CDs. Ginny owns 5 more CDs than Betty and 8 fewer than Allison. Which of the following represents the total number of CDs the three girls own? a) 3x - 3 b) 3x + 3 c) 3x + 13 d) 3x + 18

OpenStudy (anonymous):

C

OpenStudy (anonymous):

It's D. x + (x+5) + (x+13) = 3x + 18. Since better has 8 fewer CDs than Allison, and Allison has 5 more CDs than Betty, then Allison has 13 CDs.

OpenStudy (anonymous):

I give up, @pylauzier Lol

OpenStudy (anonymous):

Lol :(

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

By the way I meant to say 13 CDs more than Betty, not 13 CDs in total. That's where the x + 13 comes from.

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

Read the question, x represents the number of CDs betty own, you solved it as if x was Gina's CDs.

OpenStudy (anonymous):

Yep, I missread it!

OpenStudy (anonymous):

Thanks for pointing it out :)

OpenStudy (anonymous):

Ahh ok :D

OpenStudy (anonymous):

G = 5 + B G = A - 8 -> A = 5 + B + 8 = B + 13 B + G + A => B + ( 5 +B) +( B + 13) => 3B +18